the position x of a particle moving along xaxis at time t is given by equation t=√x +2.find the work done by the force in first 4seconds

Grade:11

5 Answers

DR STRANGE
7 years ago
Hi Vaibhav,
Again it is calculus
t=\sqrt{x}+2
x=(t-2)^{2}
after 4seconds,        S =(4-2)2=4
v=\frac{dx}{dt} =\frac{d(t-2)^{2}}{dt}=2(t-2)
a=\frac{dv}{dt}=\frac{d(2(t-2))}{dt}=2
 
W      =   F*s   =    m*a*s     =   m*2*4   =8m (you have not given mass of th particle)
 
Plz approve If you are satisfied :)
(previous question also :P)
Let me know if you still have doubts..........
 
Vaibhav
7 years ago
But mr. strange mass is not given in the question and the right answer is 0 J,,,, i dont know how but its d right answer,,,, and please view the prev. question and help me out in that too please.... and thnx for ...answering mine doubts... ;)
DR STRANGE
7 years ago
sry I mad
v=2(t-2)
initial velocity ( at time t =0)   u = 2(0-2) = -4
s = u*t + 1/2*a*t2 
after 4seconds , s= -4*4 + 1/2*2*42 = -16 + 16 = 0
s=0, hence NO work done , W=0J
And I have cleared your doubt of previous que plz check :)
Yokesh K.K.
5 years ago
Here , we have been given an equation t=√x + 2.
Let's find X. √X = t-2
X = t^2-4t+4
Now let's find velocity.
dx/dt = 2t-4.
Put dx/dt = 0.
By this way , we get the time at which the particle's velocity is zero. We get t =2 sec.
Now in the question We've been asked to find the work done in t= 4 seconds . THIS CLEARLY MEANS THAT THE PARTICLE IS REVERSING ITS DIRECTION. Hence at 4 seconds the displacement of the particle is zero . Hence work done is zero.
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

v=2(t-2)
initial velocity ( at time t =0) u = 2(0-2) = -4
s = u*t + 1/2*a*t2
after 4seconds , s= -4*4 + 1/2*2*42= -16 + 16 = 0
s=0, hence NO work done , W=0J

Thanks and Regards